OpenStudy (anonymous):

The first four terms of a sequence are shown below: 8, 5, 2, -1 Which of the following functions best defines this sequence? (4 points) f(1) = 8, f(n + 1) = f(n) + 3; for n ≥ 1 f(1) = 8, f(n + 1) = f(n) - 5; for n ≥ 1 f(1) = 8, f(n + 1) = f(n) + 5; for n ≥ 1 f(1) = 8, f(n + 1) = f(n) - 3; for n ≥ 1

OpenStudy (lochana):

can you tell me what f(1) is.

OpenStudy (anonymous):

I don't know

OpenStudy (lochana):

in other words first number of sequence

OpenStudy (lochana):

look at first number of your numbers group. and tell me first number

OpenStudy (anonymous):

8

OpenStudy (lochana):

yes. that is f(1) so f(1) = 8

OpenStudy (lochana):

now tell me second number, it will be f(2)

OpenStudy (anonymous):

I don't know?

OpenStudy (lochana):

f(2) = 5

OpenStudy (lochana):

get it?

OpenStudy (anonymous):

I guess

OpenStudy (lochana):

mm.. okay. can you see there is a relationship you can tell 5 = 8 -3 f(2) = f(1) -3

OpenStudy (anonymous):

Yes but what does it have to do with the equation

OpenStudy (lochana):

yes. so it works everywhere f(2) = f(1) -3 f(3) = f(2) -3 f(4) = f(3) -3 and you can say a common equation instead of saying every equation above, f(n+1) = f(n) -3

OpenStudy (anonymous):

Okay, so what would the answer be ?..

OpenStudy (lochana):

well it is last one.